what is the frequency of green light that has a wavelength of 497 nm? (c= 3.00 x 10^8 m/s)
1
Expert's answer
2020-04-02T08:56:26-0400

The frequency of the electromagnetic wave could be calculated using the following expression:


"f=\frac{c}{\lambda}=\frac{3.00*10^{8}m/s}{497*10^{-9}m}=6.04*10^{14}Hz."